# Enhancing Group Fairness in Federated Learning through Personalization

### v1

@@ 6.1 The Fair-FCA algorithm
Algorithm 1 Fair-FCA: A Fairness-aware Federated Clustering Algorithm
6.2 Fairness-accuracy tradeoff using Fair-FCA
Figure 6: Acc. vs. SP-fairness
We begin by conducting a numerical experiment on a synthetic dataset, to illustrate the ability of Fair-FCA to strike a balance between fairness and accuracy. We consider a total of 8 clients that could (potentially) be clustered into two clusters. Among these, 6 clients (Client ID: 2,4,5,6,7,8) have similar data distributions, with 4 clients (Client ID: 4,6,7,8) sharing identical distributions across the two protected groups $a,b$ (low fairness gap). The remaining 2 clients (Client ID: 1,3) have different data distributions compared to the first six, but they also share identical distributions across the two protected groups. We consider Fair-FCA on this problem, with $f=\texttt{SP}$, and $\gamma\in\{0,0.3,0.5,0.8,1\}$.
Figure 6 shows that when $\gamma=1$, Fair-FCA prioritizes accuracy; by design, this is attained by grouping the 6 clients having similar data distributions together ({1,3} and {2,4,5,6,7,8}). Similarly, when $\gamma=0$, Fair-FCA focuses only on SP fairness, this time clustering clients that have identical distributions on the two protected groups together ({2,5} and {1,3,4,6,7,8}). Lastly, by setting $\gamma\in(0,1)$, we can effectively account for both accuracy and SP fairness when clustering: when $\gamma=0.3$, the clusters are {2,4,5} and {1,3,6,7,8}; when $\gamma=0.5$, the clusters are {2,4,5,6} and {1,3,7,8}; and when $\gamma=0.8$, the clusters are {2,4,5,6,7} and {1,3,8}.
6.3 Numerical experiments: comparing Fair-FCA vs. IFCA (Clustered-FL)
We now compare the Clustered-FL (IFCA) algorithm of Ghosh et al. (2020) and our Fair-FCA algorithm with $\gamma=0.5$ and with $f=\texttt{SP}$, in terms of local accuracy and local group fairness. We proceed with the same experiment setting as in Section 4. (We also provide experiment results using the original Retiring Adult dataset without feature scaling in Appendix E.5, and experiments with EqOp as the fairness metric in Appendix E.3.)
Figure 7 verifies that under different datasets (ACSEmployment, ACSIncome) and different protected attributes (race, sex), Fair-FCA exhibits a fairness-accuracy tradeoff (improved fairness performance at the expense of accuracy degradation) consistent with the choice of $\gamma=0.5$ compared to the Clustered-FL algorithm.
(a) Employment (Race)
(b) Employment (Sex)
(c) Income (Sex)
Figure 7: Fair-FCA with different tasks and protected attributes ($f=\texttt{SP}$)

### v2

@@ 5.2 Fairness-accuracy tradeoff using Fair-FCA and Fair-FL+HC
Algorithm 1 Fair-FCA
Input: Number of rounds $k_{1}$/$k_{2}$ before/after clustering, number of clients $n$, number of local updates $E$, learning rate $\eta$, Model initialization $\theta$, Set of hyperparameters for the hierarchical clustering algorithm $P$, fairness-accuracy tradeoff $\gamma$, fairness $f\in\{\texttt{SP},\texttt{EqOp},\texttt{EO}\}$.
for each round $t\in[k_{1}]$ do
       Output FedAvg model: $\theta=\texttt{FedAvg}(\theta,n)$
end for
for each client $i\in[n]$ in parallel do
       Initialize $\theta_{i}=\theta$
       Perform $E$ steps of local update: $\theta_{i}=\theta_{i}-\eta\nabla_{\theta_{i}}F(Z_{i},\theta_{i})$
end for
[…]
for each cluster $c\in\text{Cluster}$ in parallel do
       for each client $i\in c$ in parallel do
             Perform FedAvg procedures: $\texttt{FedAvg}(\theta_{i},n_{c}),n_{c}\in c$
       end for
end for
Algorithm 2 Fair-FL+HC